OpenStudy (anonymous):

Let f(x) = 3x – 42 and g(x) = 5x. What is (f ◦ g)(x) ? Use complete sentences to describe the method you used to solve this problem.

OpenStudy (anonymous):

f(g(x)) = f(5x) let's start there... (not the answer)

OpenStudy (anonymous):

than what

OpenStudy (anonymous):

look at f(x) and its rule... f(x) = 3 * x - 42: f(7) = 3 * 7 - 42 f(23) = 3 * 23 - 42 f(pi) = 3 * pi -42 f(mn) = 3 * (mn) - 42 what about f(5x) = ???